Find the perimeter of a rectangle if its length measures 17 centimeters and its width measures 4 centimeters. Include correct units with your solution.
Answer by checkley77(12844)

P=2L+2W
P=2*17+2*4
P=34+8=42 CENTIMETERS IS THE PERIMETER.